Calculate the internal energy of 1 mole of an ideal gas at 250ºC.

Grade:11

1 Answers

Navjyot Kalra
8 years ago
To find out the internal energy of the ideal gas, substitute 1 mole for n, 250̊ C for temperature T and 8.31 J/ mol. K for the gas constant R in the equation E = 3/2 nRT,
E = 3/2 nRT
= 3/2 (1 mol) (8.31 J/ mol. K) (250̊ C)
= (1.5)( 1 mol) (8.31 J/ mol. K) (250+273) K
= (1.5)( 1 mol) (8.31 J/ mol. K) (523) K
= 6519.195 J
Rounding off to three significant figures, the internal energy of the ideal gas will be 6520 J.